Two planes leave Chicago - one traveling due east and the other due west. The first travels 100 miles per hour faster than the second. How long will it be until they are 2000 miles apart? ( If there is not enough information given, indicate that no solution can be found)

Answer by checkley77(12569)

D=RT
2,000=(100+X+100)T
2,000=(200+X)T
T=2,000/(200+X)
NO UNIQUE SOLUTION WHEN YOU HAVE 2 UNKNOWNS IN 1 EQUATION.
EX:
T=2,000/(200+50)
T=2,000/250
T=8 HOURS.
T=2,000/(200+100)
T=2,000/300
T=6.667 HOURS.
T=2,000/(200+200)
T=2,000/400
T=5 HOURS.